LINEAR ALGEBRA

2. (6 marks) Find the best line y=c+dt to fit y=1, 1, 2, 2 at times t=-1, 0, 1, 2. (Use the least squares approximation.)
Answer #1

we have the following D no number of points = 4 -1 + 0 + 1 + 2 = 2 2) Et 3) Et +1²+²+1²+2² = 6 1) Eys 1+1 +2 +2 5) Eyt = IX (